Problem 11. [4 pts] Emily DiDonato is thinking of buying Berkshire Hathaway Inc. (BRK-B) priced at $200 per share. She assumes that that risk-free rate is about 5% and the market risk premium (MRP) is 10%. If she thinks, the stock will rise to $320 per share by the end of the year, what beta would it need to have for this expectation to be consistent with the CAPM? (Show your calculations)Problem 12. 4 pts] Taylor Hill wants to invest $90,000 in a two-asset portfolio with a beta of 1.0. The stock XX has a beta of 2.0 while the stock YY has a beta of 0.75. How much will she invest ($) in each stock to maintain a portfolio beta of 1.0? (Show your calculations)
Answer #1

Problem 11.

Return in one year = (320-200)/200 = 60%

Using CAPM model,

r = rf +Beta(rm - rf)

Beta = (0.6 - 0.05)/0.1

Beta = 5.5

Problem 12.

Portfolio beta = w1 * Beta1 + w2 + Beta2

For portfolio beta = 1

1 = w1 * 2 + (1 - w1) * 0.75

So,

w1 = Weight of Stock XX = 20%

w2 = Weight of Stock YY = 80%